# Efficient Annotator Reliability Assessment with EffiARA

### v2

@@ Abstract
Data annotation is an essential component of the machine learning pipeline; it is also a costly and time-consuming process. With the introduction of transformer-based models, annotation at the document level is increasingly popular; however, there is no standard framework for structuring such tasks. The EffiARA annotation framework is, to our knowledge, the first project to support the whole annotation pipeline, from understanding the resources required for an annotation task to compiling the annotated dataset and gaining insights into the reliability of individual annotators as well as the dataset as a whole. The framework’s efficacy is supported by two previous studies: one improving classification performance through annotator-reliability-based soft label aggregation and sample weighting, and the other increasing the overall agreement among annotators through removing identifying and replacing an unreliable annotator. This work introduces the EffiARA Python package and its accompanying webtool, which provides an accessible graphical user interface for the system. We open-source the EffiARA Python package at https://github.com/MiniEggz/EffiARA and the webtool is publicly accessible at
@@
Owen Cook, Jake Vasilakes, Ian Roberts and Xingyi Song School of Computer Science, University of Sheffield Correspondence: oscook1@sheffield.ac.uk
@@ 1 Introduction
Labelled data is the foundation of training and evaluating downstream tasks in machine learning models. However, data annotation is often an expensive and time-consuming process, significantly affecting the quality of model training. Obtaining annotations from experts is ideal, but this expertise is often logistically and financially costly.
Crowd-sourcing platforms such as Amazon’s Mechanical Turk [^1] https://www.mturk.com/ and CrowdFlower (now Figure-Eight) [^2] https://www.appen.com/ai-data/data-annotation provide a cheaper alternative by using non-expert annotators; this generally results in lower quality annotations with higher levels of inter-annotator disagreement (Nowak and Rüger, 2010). Effectively collecting, evaluating and managing annotator disagreement is essential in addressing these challenges.
We introduce EffiARA (Efficient Annotator Reliability Assessment) framework supports annotation quality assessment and management throughout the annotation process, allowing users to:
@@
  • Generate labels for each annotator;
@@ 2 Related Work
Annotation Frameworks.
There have been many attempts to formalise the annotation process for a number of annotation tasks and a range of tools are available. Many frameworks focus on sequence labelling tasks such as POS tagging and named-entity recognition (Bird and Liberman, 2001; Cornolti et al., 2013; Bontcheva et al., 2013; Lin et al., 2019). More recently, with the introduction of pre-trained LLMs capable of document-level processing, document annotation tools and frameworks have been created such as GATE Teamware 2 (Wilby et al., 2023). A number of annotation frameworks are task-specific, aiming to provide a set of guidelines and tools for following them, for example event ordering (Cassidy et al., 2014), biodiversity information extraction (Lücking et al., 2022), and surgical video analysis (Meireles et al., 2021).
Annotator Agreement.
@@
The soft label approach incorporates a level of subjectivity into aggregated labels for each data point and has been shown to improve both classification performance and model calibration (Wu et al., 2023; Cook et al., 2024). Popular methods of label aggregation include majority voting (hard-label only), Dawid and Skene (1979), GLAD (Whitehill et al., 2009), and MACE (Hovy et al., 2013) for categorical data; these methods have been implemented in Python as part of the Crowd-Kit tool (Ustalov et al., 2021).
Annotator Reliability.
Assessing annotator reliability can be used to assess the quality of individual annotators and may be used to understand the quality of data available, remove low-reliability annotators (Cook et al., 2025), inform the training of machine learning models through aggregating soft labels with reliability (Dawid and Skene, 1979; Wu et al., 2023), or affect the loss function during training (Cao et al., 2023; Cook et al., 2024). There are different approaches to assessing annotator reliability, such as learning through Expectation Maximisation (Cao et al., 2023) or directly inferring the reliability of an annotator from their agreement with others (Inel et al., 2014; Dumitrache et al., 2018; Cook et al., 2024, 2025). Through impacting the generation of soft labels or directly impacting the training loss, more information about which annotations are more trustworthy is provided, leading to more performant and robust models. An alternative approach to assessing annotator reliability involves comparing annotators’ annotations to a set of gold-standard labels (Barthet et al., 2023); this approach is often used to filter out bad annotators. All three approaches have been shown to improve model performance on classification tasks when compared to methods that trust each annotator equally.
@@ 3 EffiARA Python Package
The EffiARA annotation framework structures the annotation process from start-to-finish. It distributes samples to annotators, generates and aggregates labels, computes inter- and intra-annotator agreement, and assesses annotator reliability. A visual representation of the EffiARA pipeline is provided in Figure 1 and we describe each stage in detail below.
@@
The package relies on a number of core Python libraries. Two fundamental libraries required by the EffiARA framework are NumPy (Oliphant et al., 2006; Harris et al., 2020) and pandas (McKinney et al., 2011) for efficient mathematical operations on arrays and the manipulation of data.
@@ 3.1 Sample Distribution
The first stage in the EffiARA pipeline enables annotation coordinators to estimate resource requirements: how many annotators are needed, how much time is required from each annotator, and how many samples can be produced given the time and number of annotators. Once resources have been finalised, data points can be distributed among annotators with the EffiARA distribution algorithm, which ensures annotator agreement can be effectively assessed (Cook et al., 2024).
@@ 3.2 Data Annotation
The sample allocations obtained in the previous step can then be used to assign samples to annotators and complete the annotation process using existing tools such as GATE Teamware 2 or Amazon’s Mechanical Turk.
@@ 3.3 Label Generation
Label generation involves transforming raw annotations obtained from annotators into numeric encodings compatible with annotator agreement metrics (such as Cohen’s Kappa, Fleiss’ Kappa, Krippendorff’s alpha, or cosine similarity) and model training. These transformations may be at the individual annotator level (for example, transforming first- and second-choice annotations into a categorical distribution) or at the data point level (aggregating annotations from multiple annotators).
@@
For annotator agreement calculations, only add_annotation_prob_labels must be implemented. To instantiate a class inheriting from LabelGenerator, the user must provide a list of annotator names and the label mapping (a dictionary where the key is the value represented in the DataFrame and the value is a numeric representation). This enables the extraction of individual annotations and their representation as a distribution across the available classes.
We provide a number of preset label generators, such as: the DefaultLabelGenerator, for the cases in which no special label aggregation is necessary; the EffiLabelGenerator, mirroring the label generation and aggregation shown in Cook et al. (2024); the TopicLabelGenerator, for multi-label tasks such as topic-extraction (Cook et al., 2025); and the OrdinalLabelGenerator, used for ordinal annotation tasks where a number of features are labelled on a scale. With the label_generator.from_annotations method, the specific class inheriting from LabelGenerator is instantiated from the raw annotations, requiring no additional coding from the user.
@@ 3.4 Annotator Agreement & Reliability
Once labels for each annotation have been generated, inter- and intra-annotator agreement are calculated using equations introduced in Cook et al. (2024). Annotator agreement can then be visualised in a 2D or interactive 3D graph, where each node represents an annotator and edges between annotators represent the pairwise agreement between two annotators, with the value next to each node representing an annotator’s agreement with themself. For cases with many annotators, where a graph could be unwieldy, we also provide a heatmap visualisation, where annotators are ordered by reliability; intra-annotator agreement is displayed on the diagonal. Examples of these visualisations are given in Figure 2.
@@ B Heatmap
Using these agreement calculations, annotator reliability can then be calculated, using a combination of an annotator’s intra-annotator agreement and average inter-annotator agreement, weighted by an $\alpha$ parameter controlling the strength of intra-annotator agreement from 0-1. The resulting agreement values are centered around 1, enabling the recursive inter-annotator agreement calculation from Cook et al. (2024). The reliability values can then be accessed and utilised, potentially removing certain annotators from the annotation process (Cook et al., 2025). They may also be utilised in label aggregation (in a LabelGenerator) or used to weight the loss function in model training (Cook et al., 2024).
@@
On instantiating an Annotations class, the annotator graph (supported by the NetworkX library (Hagberg et al., 2008)) is initialised with each annotator equally reliable. Intra-annotator agreement is first calculated for each annotator node with the calculate_intra_annotator_agreement instance method, using samples each user has annotated twice themselves. Inter-annotator agreement is then calculated between each user, utilising the overlap_threshold to decide whether there is sufficient overlap between the two annotators to assess agreement. Here, the pairwise_agreement function is used as a common interface to interface with the implemented pairwise agreement metrics in the agreement module. Python modules used to handle agreement calculations include the Krippendorff library (Castro, 2017) for Kippendorff’s alpha and Scikit-Learn (Pedregosa et al., 2011) for Cohen’s Kappa and Fleiss’ Kappa. NumPy and pandas are also used for vector calculations and manipulation of the data to obtain pair annotations.
@@
Inter- and intra-annotator agreement values can also be easily accessed via the graph itself using the NetworkX API and the __getitem__ method of the Annotations class. The graph and heatmap agreement visualisations shown in Figure 2 utilise Matplotlib (Tosi, 2009) and Seaborn (Waskom, 2021) and are displayed using the display_annotator_graph and display_agreement_heatmap methods, respectively.
@@ 4 EffiARA Webtool
The webtool supports common tasks within the annotation pipeline (excluding the annotation step itself). Finer-grained control and more advanced functionality may be achieved with the Python package, particularly through customisation of modules like the LabelGenerator. As the project is open-sourced, technical users are able to make their own modifications and run them as a local web-application or make a pull request to add their additional use-cases. Webtool source code available at https://github.com/MiniEggz/EffiARA-webtool.
@@ 5.1 Case Studies
In the work introducing the EffiARA annotation framework (Cook et al., 2024), utilising reliability scores in the label generation and model training stages was shown to improve classification performance. Applying a soft-label approach, using TwHIN-BERT-Large, assessing reliability with inter-annotator agreement only, intra-annotator agreement only, and a combination of both all improved classification performance. Classification performance increased an F1-macro score of 0.691 to 0.740 using the EffiARA reliability scores calculated using a reliability alpha of $0.5$.

### v3

@@ Abstract
Data annotation is an essential component of the machine learning pipeline; it is also a costly and time-consuming process. With the introduction of transformer-based models, annotation at the document level is increasingly popular; however, there is no standard framework for structuring such tasks. The EffiARA annotation framework is, to our knowledge, the first project to support the whole annotation pipeline, from understanding the resources required for an annotation task to compiling the annotated dataset and gaining insights into the reliability of individual annotators as well as the dataset as a whole. The framework’s efficacy is supported by two previous studies: one improving classification performance through annotator-reliability-based soft-label aggregation and sample weighting, and the other increasing the overall agreement among annotators through removing identifying and replacing an unreliable annotator. This work introduces the EffiARA Python package and its accompanying webtool, which provides an accessible graphical user interface for the system. We open-source the EffiARA Python package at https://github.com/MiniEggz/EffiARA and the webtool is publicly accessible at
@@
Owen Cook, Jake Vasilakes, Ian Roberts and Xingyi Song School of Computer Science, University of Sheffield Correspondence: {oscook1, x.song}@sheffield.ac.uk
@@ 1 Introduction
Labelled data is the foundation of model training and evaluating downstream tasks in machine learning models. However, data annotation is often an expensive and time-consuming process, significantly affecting the quality of model training. Obtaining annotations from experts is ideal, but this expertise is often logistically and financially costly.
Crowd-sourcing platforms such as Amazon’s Mechanical Turk [^1] https://www.mturk.com/ and CrowdFlower (now Figure-Eight) [^2] https://www.appen.com/ai-data/data-annotation provide a cheaper alternative by using non-expert annotators; this generally results in lower quality annotations with higher levels of inter-annotator disagreement (Nowak and Rüger, 2010). Effectively collecting, evaluating and managing annotator disagreement is essential in addressing challenges of data quality.
We introduce the EffiARA (Efficient Annotator Reliability Assessment) framework, which supports annotation quality assessment and management throughout the annotation process, allowing users to:
@@
  • Generate labels from the annotations of each annotator;
@@ 2 Related Work
2.1 Annotation Frameworks
There have been many attempts to formalise the annotation process for a number of annotation tasks and a range of tools are available. Many frameworks focus on sequence-labelling tasks such as POS tagging and named-entity recognition (Bird and Liberman, 2001; Cornolti et al., 2013; Bontcheva et al., 2013; Lin et al., 2019). More recently, with the introduction of pre-trained LLMs capable of document-level processing, document annotation tools and frameworks have been created, such as GATE Teamware 2 (Wilby et al., 2023). INCEpTION (Klie et al., 2018; De Castilho et al., 2024) handles various aspects of annotator management, including workload distribution, as well as the annotation itself, with a customisable UI, supporting span and document annotation. It also enables active learning but does not aim to model annotator or dataset reliability explicitly. A number of annotation frameworks are task-specific, aiming to provide a set of guidelines and tools for following them, for example event ordering (Cassidy et al., 2014), biodiversity information extraction (Lücking et al., 2022), and surgical video analysis (Meireles et al., 2021).
2.2 Annotator Agreement
@@ 2.3 Annotation Aggregation
The soft-label approach incorporates a level of subjectivity into aggregated labels for each data point and has been shown to improve both classification performance and model calibration (Wu et al., 2023; Cook et al., 2024). Popular methods of label aggregation include majority voting (hard-label only), Dawid and Skene (1979), GLAD (Whitehill et al., 2009), and MACE (Hovy et al., 2013) for categorical data; these methods have been implemented in Python as part of the Crowd-Kit tool (Ustalov et al., 2021).
2.4 Annotator Reliability
Assessing annotator reliability can be used to assess the quality of individual annotators and may be used to understand the quality of data available, remove low-reliability annotators (Cook et al., 2025), inform the training of machine learning models through aggregating soft labels with reliability (Dawid and Skene, 1979; Wu et al., 2023), or affect the loss function during training (Cao et al., 2023; Cook et al., 2024). There are different approaches to assessing annotator reliability, such as learning through Expectation Maximisation (Cao et al., 2023) or directly inferring the reliability of an annotator from their agreement with others (Inel et al., 2014; Dumitrache et al., 2018; Cook et al., 2024, 2025). Through impacting the generation of soft labels, or directly impacting the training loss, more information about which annotations are more trustworthy is provided, leading to more performant and robust models. An alternative approach to assessing annotator reliability involves comparing annotators’ annotations to a set of gold-standard labels (Barthet et al., 2023); this approach is often used to filter out bad annotators. All three approaches have been shown to improve model performance on classification tasks when compared to methods that trust each annotator equally.
@@ 3 EffiARA Python Package
The EffiARA annotation framework structures the annotation process from start to finish. It distributes samples to annotators, generates and aggregates labels, computes inter- and intra-annotator agreement, and assesses annotator reliability. A visual representation of the EffiARA pipeline is provided in Figure 1 and we describe each stage in detail below. The framework has been implemented in Python due to its extensive use in NLP research, increasing its ease of use and integration.
@@
The package relies on a number of core Python libraries. Two fundamental libraries required by the EffiARA framework are NumPy (Oliphant et al., 2006; Harris et al., 2020) and pandas (McKinney et al., 2011), used for efficient mathematical operations on arrays and the manipulation of data.
@@ 3.1 Sample Distribution
The first stage in the EffiARA pipeline enables annotation coordinators to estimate resource requirements: how many annotators are needed, how much time is required from each annotator, and how many samples can be produced, given the time and number of annotators. Once resources have been finalised, data points can be distributed among annotators with the EffiARA distribution algorithm, which ensures annotator agreement can be effectively assessed (Cook et al., 2024).
@@ 3.2 Data Annotation
The sample allocations obtained in the previous step can then be used to assign samples to annotators and complete the annotation process using existing tools such as GATE Teamware 2 (Wilby et al., 2023) or Amazon’s Mechanical Turk.
@@ 3.3 Label Generation
Label generation involves transforming raw annotations obtained from annotators into numeric encodings compatible with annotator agreement metrics (such as Cohen’s Kappa, Fleiss’ Kappa, Krippendorff’s alpha, or cosine similarity) and model training. These transformations may be at the individual annotator level (for example, transforming first- and second-choice annotations into a categorical distribution), or at the data point level (aggregating annotations from multiple annotators).
@@
For annotator agreement calculations, only add_annotation_prob_labels must be implemented. To instantiate a class inheriting from LabelGenerator, the user must provide a list of annotator names and the label mapping: a dictionary where the key is the value represented in the DataFrame and the value is a numeric representation. This enables the extraction of individual annotations and their representation as a distribution across the available classes.
We provide a number of preset label generators: the DefaultLabelGenerator, for the cases in which no special label aggregation is necessary; the EffiLabelGenerator, mirroring the label generation and aggregation shown in Cook et al. (2024); the TopicLabelGenerator, for multi-label tasks such as topic-extraction (Cook et al., 2025); and the OrdinalLabelGenerator, used for ordinal annotation tasks where a number of features are labelled on a scale. With the label_generator.from_annotations method, the specific class inheriting from LabelGenerator is instantiated from the raw annotations, requiring no additional coding from the user.
@@ 3.4 Annotator Agreement & Reliability
Once labels for each annotation have been generated, inter- and intra-annotator agreement are calculated using equations introduced in Cook et al. (2024). Annotator agreement can then be visualised in a 2D or interactive 3D graph, where each node represents an annotator and edges between annotators represent the pairwise agreement between two annotators, with the value next to each node representing an annotator’s agreement with themself. For cases with many annotators, where a graph could be unwieldy, we also provide a heatmap visualisation, where annotators are ordered by reliability; note that intra-annotator agreement is displayed on the diagonal. Examples of these visualisations are given in Figure 2.
@@ B Heatmap
Using these agreement values, annotator reliability can then be calculated, using a combination of an annotator’s intra-annotator agreement and average inter-annotator agreement, weighted by an $\alpha$ parameter controlling the strength of intra-annotator agreement from 0 to 1. The resulting agreement values are centered around 1, enabling the recursive inter-annotator agreement calculation from Cook et al. (2024). The reliability values can then be accessed and utilised, potentially removing certain annotators from the annotation process (Cook et al., 2025). Reliability scores may also be utilised in label aggregation (in a LabelGenerator) or used to weight the loss function in model training (Cook et al., 2024).
@@
On instantiating an Annotations class, the annotator graph (supported by the NetworkX library (Hagberg et al., 2008)) is initialised with each annotator equally reliable. Intra-annotator agreement is first calculated for each annotator node with the calculate_intra_annotator_agreement instance method, using data points each user has annotated twice themselves. Inter-annotator agreement is then calculated between each user, utilising the overlap_threshold to decide whether there is sufficient overlap between the two annotators to assess agreement. Here, the pairwise_agreement function is used as a common interface to the implemented pairwise agreement metrics in the agreement module. Python modules used to handle agreement calculations include the Krippendorff library (Castro, 2017) for Krippendorff’s alpha and Scikit-Learn (Pedregosa et al., 2011) for Cohen’s Kappa and Fleiss’ Kappa. NumPy and pandas are also used for vector calculations and manipulation of the data to obtain pair annotations.
@@
Inter- and intra-annotator agreement values can also be easily accessed via the graph itself using the NetworkX API and the __getitem__ method of the Annotations class. The graph and heatmap agreement visualisations shown in Figure 2 utilise Matplotlib (Tosi, 2009) and Seaborn (Waskom, 2021), and they are displayed using the display_annotator_graph and display_agreement_heatmap methods respectively.
@@ 4 EffiARA Webtool
The webtool supports common tasks within the annotation pipeline (excluding the annotation step itself). Finer-grained control and more advanced functionality can be achieved with the Python package, particularly through customisation of modules like the LabelGenerator. As the project is open-sourced, technical users are able to make their own modifications and run them as a local web-application or make a pull request to add their additional use-cases. The webtool source code is available at https://github.com/MiniEggz/EffiARA-webtool.
@@ 5.1 Case Studies
In the work introducing the EffiARA annotation framework (Cook et al., 2024), utilising reliability scores in the label generation and model training stages was shown to improve classification performance. Applying a soft-label approach, using TwHIN-BERT-Large, assessing reliability with inter-annotator agreement only, intra-annotator agreement only, and a combination of both all improved classification performance. Classification performance increased from an F1-macro score of 0.691 to 0.740 using the EffiARA reliability scores calculated using a reliability alpha of $0.5$. The dataset used in this study was of low-to-moderate agreement, highlighting the framework’s utility in datasets containing disagreement.
@@ 6 Conclusion and Future Work
7 Limitations
EffiARA has been developed primarily to support document classification annotation, with the flexibility to accommodate other annotation types. While the design accounts for such flexibility, span annotation functionality, for example, has not yet been implemented and would require technical expertise to integrate. One key future addition may include the integration of Krippendorff’s unitising alpha (Krippendorff et al., 2016) for spans.
Some annotation tasks may require task-specific label generators. Although the framework includes preset label generators, less common tasks may require further customisation. This may present a usability barrier for non-technical users. We aim to address this through future development and community contributions of further task-specific components.
While EffiARA has been tested in situations with disagreement, it has been assumed that there is a ground-truth label for each data point. Future work will investigate the extension of EffiARA to tasks where there may not be a single ground-truth label, but potentially multiple subjective, and equally valid, true labels.
@@ 8 Ethical Impact
9 Acknowledgements
This work was supported by the UK’s innovation agency (Innovate UK) grant 10039055 (approved under the Horizon Europe Programme as vera.ai, EU grant agreement 101070093) vera.ai. JV is supported by the European Media and Information Fund (EMIF) managed by the Calouste Gulbenkian Foundation [^3] The sole responsibility for any content supported by the European Media and Information Fund lies with the author(s) and it may not necessarily reflect the positions of the EMIF and the Fund Partners, the Calouste Gulbenkian Foundation and the European University Institute. under the “Supporting Research into Media, Disinformation and Information Literacy Across Europe” call (ExU – project number: 291191).444exuproject.sites.sheffield.ac.uk We also acknowledge the IT Services at The University of Sheffield, utilising their High Performance Computing clusters in both of our case studies.